Show that simple harmonic motion y(t) = C1 cos ωt + C2 sin ωt can be written as:


a. y(t) = A sin(ωt + ϕ0)


b. y(t) = A cos(ωt + ϕ1)


1
Expert's answer
2022-05-02T18:06:11-0400

"A=\sqrt{C_1^2+C_2^2},"

"\phi_0=\arccos\frac{C_2}{\sqrt{C_1^2+C_2^2}},"

"\phi_1=-\arcsin\frac{C_2}{\sqrt{C_1^2+C_2^2}}."